OpenStudy (erniellyd28):

y equals 2x minus 3 (this type of problwem has to do with graphing)

jimthompson5910 (jim_thompson5910):

y = 2x - 3 is in the form y = mx + b where m = 2 and b = -3

jimthompson5910 (jim_thompson5910):

so the slope is 2 or 2/1 and the y-intercept is the point (0, -3)

jimthompson5910 (jim_thompson5910):

to graph y = 2x - 3, you start with the y-intercept (0, -3) then you go up 2 units and over to the right 1 unit (this is from the slope rise/run = 2/1) this will make you go from (0,-3) to (1, -1)

jimthompson5910 (jim_thompson5910):

now that you have the points (0,-3) and (1, -1), you just draw a straight line through them to graph y = 2x - 3

OpenStudy (erniellyd28):

thank you so very much for all your help i really appreciate it

OpenStudy (erniellyd28):

before i let u go may i ask for your help on one more problem please

jimthompson5910 (jim_thompson5910):

sure thing, what is it

OpenStudy (erniellyd28):

find the xminus and yminus intercepts 3x plus 4y equals 12

jimthompson5910 (jim_thompson5910):

to find the x-intercept, you plug in y = 0 and solve for x

jimthompson5910 (jim_thompson5910):

3x + 4y = 12 3x + 4(0) = 12 3x + 0 = 12 3x = 12 x = 12/3 x = 4

jimthompson5910 (jim_thompson5910):

so the x-intercept is the point (4, 0)

jimthompson5910 (jim_thompson5910):

to find the y-intercept, you plug in x = 0 and solve for y I'll let you do this one. Tell me what you get

OpenStudy (erniellyd28):

i am totally lost im sorry

jimthompson5910 (jim_thompson5910):

do you see how I got the x-intercept?

OpenStudy (erniellyd28):

yes but i think i still got confuse

OpenStudy (erniellyd28):

im very sorry

jimthompson5910 (jim_thompson5910):

where are you stuck?

jimthompson5910 (jim_thompson5910):

don't be sorry, it's ok to not get it and ask questions